OpenStudy (cookiemonster18):

1.Find the slope of the linear line between the points (-4, 5) and (6, 1). 2.Find the slope of the linear line between the points (10, -4) and (4, -4).

Nnesha (nnesha):

\[\large\rm \frac{ y_2 -y_1 }{ x_2-x_1 }\] formula to find slope between two points \[\rm (x_1,y_1)(x_2,y_2)\]

Nnesha (nnesha):

\[\rm \large \rm (\color{Red}{x_1},\color{blue}{y_1})(\color{red}{x_2},\color{blue}{y_2} )\]\[\rm \large \rm (\color{Red}{-4},\color{blue}{5})(\color{red}{6},\color{blue}{1} )\] plugin x's and y's values into the formula then simplify

OpenStudy (cookiemonster18):

x=-10 and y=4??? help me :( @Nnesha

Nnesha (nnesha):

`1.Find the slope of the linear line between the points (-4, 5) and (6, 1).` first question

Nnesha (nnesha):

(x,y) ordered pair where first number represent x-coordinate and 2nd one y-coordinate so (-4,5) where x=-4 and y=5

OpenStudy (cookiemonster18):

so x=-4 - x=6 y=5 - y=1

Nnesha (nnesha):

yes correct so now plug that into the formula (x_1,y_1)(x_2,y_2) x_1 means first x y_2 means 2nd y

Nnesha (nnesha):

but remember y values should be at the numerator

OpenStudy (cookiemonster18):

x=-4 - x=6 y=5 - y=1 equals x=2?? y=4??

Nnesha (nnesha):

hmm if x=-4 and x=6 then what would be -4-6 = ?

Nnesha (nnesha):

\[\large\rm \frac{ y_2 -y_1 }{ x_2-x_1 }\] this is the formula plugin x =-4 and x=6 \[\large\rm \frac{ y_2 -y_1 }{ 6-(-4) }\] same with y values

OpenStudy (cookiemonster18):

-4-6=-10 6-(-4)=10

Nnesha (nnesha):

yes correct so so what would be the final answer ?\[\frac{ 1-5 }{ 6-(-4) } =\frac{-4}{10}\] you will get the same answer doens't matter if you write y_1 first or y_2

Nnesha (nnesha):

now just reduce the fraction

OpenStudy (cookiemonster18):

-4/10=-0.4

Nnesha (nnesha):

slope = rise over run so keep that in fraction form divide top and bottom of the fraction by 2

OpenStudy (cookiemonster18):

-4/10=2/5 am I correct?

Nnesha (nnesha):

ye but it should be negative bec negative divide by positive = negative

OpenStudy (cookiemonster18):

sorry I had to fix it and I guess I may have deleted the negative sign so the answer is -2/5

Nnesha (nnesha):

correct.

OpenStudy (cookiemonster18):

thank you can you tell me if im correct on #2?

Nnesha (nnesha):

sure try it first show ur work i'll check :=))

OpenStudy (cookiemonster18):

x=10,4 y=-4,-4

Nnesha (nnesha):

ye those are x and y values not plug that into the formula

Nnesha (nnesha):

now**

OpenStudy (cookiemonster18):

10-4=6 -4(-4)=16 6/16=3/8 am I correct the answer is 3/8???

Nnesha (nnesha):

remember the formula is in other words \[\rm \frac{ change~ in~y's }{ change ~ in~x's }\]

Nnesha (nnesha):

subtract y values in the numerator and subtract x's values in the denomiantor so we should subtract them not multiply

OpenStudy (cookiemonster18):

-4(-4)=16 10-4=6 16/6=8/3???

Nnesha (nnesha):

don't do it serrate use the formula again we are not multiplying y values we should subtract them

Nnesha (nnesha):

\[\large\rm \frac{ y_2 -y_1 }{ x_2-x_1 }\] write like this in fraction form

Nnesha (nnesha):

separate **

OpenStudy (cookiemonster18):

-4(-4)=16 4-10=-6 16/-6???

Nnesha (nnesha):

like i said before we are NOT multiplying

Nnesha (nnesha):

subtract y values not multiply 4(-4) isn't correct

OpenStudy (cookiemonster18):

it should be 2 negatives -4-4=-8

Nnesha (nnesha):

no that's the reason i'm saying use the formula

Nnesha (nnesha):

there is a negative sign in the formula here \[\large\rm \frac{ y_2 -y_1 }{ x_2-x_1 }\] substitute y_2 for -4 and y_1 for -4

OpenStudy (cookiemonster18):

as in y_1 you mean 10 and y_2 you mean 4

Nnesha (nnesha):

10 and 4 are x-coordinate not y -coordinate x_1 =10 x_2=4 y_1=-4 y_2=-4

Nnesha (nnesha):

 substitute x's and y's for the values